OpenStudy (kakarot):

-10u = -9u + 8 What is the value of u?

OpenStudy (idealist10):

-10u=-9u+8 -10u-(-9u)=8 -10u+9u=8 -u=8 u=-8 Answer: -8

OpenStudy (anonymous):

u=-8

OpenStudy (kakarot):

Yep that was it thanks

OpenStudy (idealist10):

You're welcome!